OpenStudy (anonymous):

Tell whether the lines for each pair of equations are parallel, perpendicular, or neither.y- 3 = 8/3(x+2); (-2,3)

OpenStudy (anonymous):

Parallel lines have equal slopes; perpendicular lines have opposite reciprocal slopes.

OpenStudy (anonymous):

(And by "opposite", I mean negative).

OpenStudy (anonymous):

yes I know that. I juat need to see someone else work it, so I can compare answers.

OpenStudy (anonymous):

Well, you only give one equation, so there's nothing we can do at this point.

OpenStudy (anonymous):

It just requires an equation.

OpenStudy (anonymous):

Using the formula: y - y1 = m(x - x1)

OpenStudy (anonymous):

Thanks for your attempt in answering yakeylee! I got the answer, it was y - 3 = (3/8(x + 2)

OpenStudy (anonymous):

For your effort I will reward you :)